OpenStudy (samigupta8):

Let f(x)=√log(x^2+kx+k+1) /x^2+k If f(x) is continuous for all x€R ,then the range of k is

OpenStudy (samigupta8):

The surd is just over numerator and not over denominator term

OpenStudy (owlcoffee):

For this function to be continous we just have to analyze the denominator and the function inside the logarithm. So, first of all, we will create the following body of equations: \[x^2+kx+k+1 > 0\] \[x^2+k \ne 0\]

OpenStudy (samigupta8):

Done! I did the same ..

OpenStudy (anonymous):

it is a guess only. \[x^2+kx+\frac{ k^2 }{ 4 }-\frac{ k^2 }{ 4 }+k+1>0\] \[\left( x+\frac{ k }{ 2 } \right)^2>\frac{ k^2 }{ 4 }-k-1\] \[\left( x+\frac{ k }{ 2 } \right)^2>\left( \frac{ k^2-4k+4-4 }{ 4 } \right)-1\] \[\left( x+\frac{ k }{ 2 } \right)^2>\left( \frac{ k-2 }{ 2 } \right)^2-2>-2\] i don't know if i am correct.

OpenStudy (samigupta8):

Alright! Bt what shall we do with this result??

OpenStudy (samigupta8):

@michele_laino

OpenStudy (michele_laino):

is your function, like this: \[f\left( x \right) = \sqrt {\log \left( {\frac{{{x^2} + kx + k}}{{{x^2} + k}}} \right)} \]

OpenStudy (samigupta8):

No , sir The surd is just over numerator

OpenStudy (samigupta8):

X^2+kx+k+1 is numerator and it is under surd sign...

OpenStudy (michele_laino):

like this one? \[f\left( x \right) = \frac{{\sqrt {\log \left( {{x^2} + kx + k} \right)} }}{{{x^2} + k}}\]

OpenStudy (samigupta8):

+1 is also there in numerator

OpenStudy (samigupta8):

Under surd sign in addition with that term u wrote

OpenStudy (michele_laino):

here it it then: \[f\left( x \right) = \frac{{\sqrt {\log \left( {{x^2} + kx + k + 1} \right)} }}{{{x^2} + k}}\]

OpenStudy (samigupta8):

Ryt!!!

OpenStudy (michele_laino):

oops..here it is then...

OpenStudy (samigupta8):

Np :)

OpenStudy (michele_laino):

we have to request these conditions: \[\begin{gathered} \log \left( {{x^2} + kx + k + 1} \right) \geqslant 0 \hfill \\ \hfill \\ {x^2} + k \ne 0 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

from the second condition, we get: \(k>0\)

OpenStudy (michele_laino):

from the first condition, we get: \[{x^2} + kx + k + 1 \geqslant 1\]

OpenStudy (michele_laino):

now, if: \[{x^2} + kx + k + 1 \geqslant 1\] then \[{x^2} + kx + k \geqslant 0\]

OpenStudy (michele_laino):

that above inequality, is equivalent to this one: \[\large \left\{ {x - \left( {\frac{k}{2} + \sqrt {\frac{{{k^2}}}{4} - k} } \right)} \right\}\left\{ {x - \left( {\frac{k}{2} - \sqrt {\frac{{{k^2}}}{4} - k} } \right)} \right\} \geqslant 0\]

OpenStudy (michele_laino):

oops.. I made a typo: \[\large \left\{ {x - \left( {\frac{{ - k}}{2} + \sqrt {\frac{{{k^2}}}{4} - k} } \right)} \right\}\left\{ {x - \left( {\frac{{ - k}}{2} - \sqrt {\frac{{{k^2}}}{4} - k} } \right)} \right\} \geqslant 0\]

OpenStudy (michele_laino):

we have the further condition: \[\frac{{{k^2}}}{4} - k \geqslant 0 \Rightarrow k \leqslant 0 \cup k \geqslant 4\] which gives \(k \geq 4\), since k has to be positive

OpenStudy (samigupta8):

Sir bt ans is different

OpenStudy (mathmale):

Look at Let f(x)=√log(x^2+kx+k+1) /(x^2+k) again. I am assuming that parentheses enclose x^2+k. If this assumption is incorrect, please take a screen shot or find some other way to communicate this function to the rest of us. Since division by zero is not defined, x^2 + k must never be zero. Note that x=0 and k = 0 is not permissible, because x^2 + k would then be zero. Next, focus on the \[\sqrt{\log (x^2+kx+k+1)}\]

OpenStudy (mathmale):

Look at Let f(x)=√log(x^2+kx+k+1) /(x^2+k) again. I am assuming that parentheses enclose x^2+k. If this assumption is incorrect, please take a screen shot or find some other way to communicate this function to the rest of us. Since division by zero is not defined, x^2 + k must never be zero. Note that x=0 and k = 0 is not permissible, because x^2 + k would then be zero. Next, focus on the \[\sqrt{\log (x^2+kx+k+1)}\]

OpenStudy (samigupta8):

Your assumption is correct parantheses has that term (x^2+k) excluded from the surd sign

OpenStudy (samigupta8):

x^2+kx+k+1>0 for log to be defined

OpenStudy (michele_laino):

I think that the subsequent conditions hold: \[\left\{ \begin{gathered} \log \left( {{x^2} + kx + k + 1} \right) \geqslant 0 \hfill \\ \hfill \\ {x^2} + kx + k + 1 > 0 \hfill \\ \hfill \\ {x^2} + k \ne 0 \hfill \\ \end{gathered} \right.\]

OpenStudy (mathmale):

Regarding your post of 5 minutes ago: Yes.. x^2+kx+k+1>0 has to be greater than 0 for that log to be defined. However, the log itself (I am not speaking of x^2+kx+k+1) could be negative, which would produce a negative argument for the sqrt function and thus would not be permissible. This is not the answer to your question, of course, but I hope this additional info to consider will be helpful.

OpenStudy (samigupta8):

Yes that would give the equation to be same as that Michele sir gave

OpenStudy (samigupta8):

Log(x^2+kx+k+1)>=0 which gives us the same equation as he gave ... X^2+kx+k+1>=1

OpenStudy (michele_laino):

we have the subsequent logic implication: \[\huge {x^2} + k \ne 0 \Rightarrow k > 0\]

OpenStudy (samigupta8):

Yes sir this part Is definitely true BT what about that k>=4 that u came up with

OpenStudy (michele_laino):

since we can write this: \[\Large \log \left( {{x^2} + kx + k + 1} \right) \geqslant 0 \Rightarrow {x^2} + kx + k + 1 \geqslant 1\]

OpenStudy (samigupta8):

Sure!

OpenStudy (michele_laino):

provided that the base of logarithm is greter than \(1\)

OpenStudy (michele_laino):

greater*

OpenStudy (samigupta8):

Well! That's not mentioned in the question though... If it's greater or less than 1

OpenStudy (michele_laino):

so, I ask, what is the base of logarithm?

OpenStudy (samigupta8):

OK sir shall I show u my work now...?

OpenStudy (michele_laino):

please wait a moment I consider the case where the base of logaritm is less than 1 and greater than zero

OpenStudy (michele_laino):

I'm doing some computations...

OpenStudy (samigupta8):

x^2+kx+k+1>=1 x^2+kx+k>=0 That means that this quadratic must have Its DETERMINANT AS <=0 So that means k^2-4k<0 That means k €[0,4]

OpenStudy (samigupta8):

Sorry open at 0

OpenStudy (michele_laino):

yes! correct!

OpenStudy (samigupta8):

I have base to be greater than 1 and most preferably it should be greater than 1 only bcoz until unless we r mentioned we have been taught that consider that case only

OpenStudy (samigupta8):

Considered*

OpenStudy (mathmale):

Michele: The base of the log here is 10; that's a given. So there's no point in wondering what would happen if the base were between 0 and 1.

OpenStudy (michele_laino):

ok! thanks! @mathmale

OpenStudy (samigupta8):

No, @mathmale anyone can wonder that becoz question never says that...

OpenStudy (samigupta8):

Base is greater than 1 or not

OpenStudy (michele_laino):

if base is less than 1, we get: \[\begin{gathered} {x^2} + kx + k + 1 \leqslant 1 \hfill \\ \hfill \\ {x^2} + kx + k \leqslant 0 \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

BT sir we can do it considering base greater than 1 becoz it's the more preferable technique and I solved the question by that and got the ans

OpenStudy (michele_laino):

more precisely: \[\begin{gathered} 0 < {x^2} + kx + k + 1 \leqslant 1 \hfill \\ \hfill \\ - 1 < {x^2} + kx + k \leqslant 0 \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

Thanks all! I got the ans ....

OpenStudy (michele_laino):

please what is the answer?